x varies inversely as y. x is 16 when y is 5.find x when y is 2.
Answer by Alan3354(69443)

x varies inversely as y. x is 16 when y is 5.find x when y is 2.
------------
x varies inversely as y --> x = k/y (k is a constant)
16 = k/5
k = 80
---
x = 80/2
x = 40
===========